OpenStudy (steve816):

This is an algebra 2 question that requires some skill, but help would be greatly appreciated!

OpenStudy (steve816):

\[\frac{ x }{ (x-1)^{2} }+\frac{ 2 }{ x }-\frac{ x+1 }{ x ^{3}-x ^{2} }\]

OpenStudy (steve816):

Simplify please!

OpenStudy (anonymous):

Looking for someone to walk you through it?

OpenStudy (steve816):

Well, actually jsut the answer to the simplified equation would be good enough!

OpenStudy (anonymous):

that's why I asked. We're not allowed to give answers here as it doesn't really benefit anyone. http://openstudy.com/code-of-conduct If you just want the answer try an online calculator like wolframalpha.com

OpenStudy (steve816):

Well then please go ahead and walk me through, that would be still great.

OpenStudy (bmmoffattb):

I'll help you.

OpenStudy (bmmoffattb):

(x/x^2-1) + (2/x) + (x+1/x^3-x^2) =(x^-1/-1) + (2/x) + (x^-4 +1)

OpenStudy (bmmoffattb):

=(x^-1/-1) + (2x^-1) + (x^-4 +1)

OpenStudy (bmmoffattb):

=(x^-1/-1) + 2x^-5 +1

OpenStudy (bmmoffattb):

Not sure what to do from there. It's been a couple years since I've done Algebra 2.

OpenStudy (bmmoffattb):

I'm sure there are laws that tell you what to do.

OpenStudy (anonymous):

Factor the third denominator to get \[x^2(x-1)\]

OpenStudy (anonymous):

So now the 3 denominators are (x- 1)², x, and x²(x -1) the least common denominator needs to have the maximum of each factor. So we need both x and (x -1) twice LCD = \(x^2(x-1)^2\)

OpenStudy (anonymous):

To make all 3 fraction match the LCD you have to multiply them by different factors. Multiply the first by \[\frac{ x^2 }{ x^2 }\] Multiply the 2nd by \[\frac{ x(x-1)^2 }{ x(x-1)^2 }\] Multiply the 3rd by \[\frac{ x-1 }{ x-1 }\]

OpenStudy (anonymous):

the problem is now\[\frac{ x(x^2) }{ x^2(x-1)^{2} }+\frac{ 2x(x-1) }{ x(x)(x-1) }-\frac{ (x+1)(x-1) }{ x^2(x-1)^2 }\]

OpenStudy (anonymous):

\[\frac{ x^2+2x(x-1)-(x+1)(x-1) }{ x^2(x-1)^2 }\] Simplify the numerator and see if anything cancels with the denominator after